Status Report

NASA Space Station On-Orbit Status 29 Jun 2004

By SpaceRef Editor
June 29, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 29 Jun 2004

All ISS systems continue to function nominally except those noted previously or below.  

The current sleep/wake cycle delayed crew wakeup time today by four hours, to 6:00am EDT, with bedtime tonight at 11:00pm.  Tomorrow, wakeup is set for 9:00am and sleep for 5:00am the next day.  By Friday, the schedule will have returned to the usual pattern.

All systems are reported ready for tomorrow’s Orlan EVA (hatch open: 5:40pm EDT). 

In a repeat of last week’s spacewalk preparations, CDR Padalka and FE/SO Fincke restored their Orlan-M suits to readiness by filling the DIDBs (disposable in-suit drink bags) and installing them in the suits.  Mike Fincke’s “skaphandr” received a fresh BK-3 oxygen bottle, and both suits were equipped with new LiOH (lithium hydroxide) cartridges for filtering carbon dioxide.

Padalka also equipped both Orlans with “Pille-MKS” radiation sensors, which he obtained from their regular exposure locations in the RS after recording their dose measurements.  A third sensor, for background readings, was placed in the SM cabin.   [“Pille” has ten sensors normally situated at various locations in the RS (port cabin window, stbd cabin window, ASU toilet facility, control panel, etc.).  Dosage values are called down or downlinked via Regul Paket/Email or OCA.]

In preparation for the EVA, which requires the station reconfigured to “unmanned mode”, Gennady powered down the new “Sputnik-SM” ham radio system in the SM.

FE Fincke got the Ku-band TV system ready for covering the EVA by reconnecting the UOP (utility outlet panel) bypass power cable to the RWS DCP (robotics workstation display & control panel), followed by configuring the Sony PD100 camcorders in the Node and Lab for situational (internal) monitoring during the EVA.

Since hatches will be closed between the Russian and U.S. segments (and also between the SM and the DC1), later tonight Mike will reset the OpsLAN onboard laptop network accordingly.

At ~2:00pm, the crew will conduct another review of the final spacewalk timeline, which included new times for the RSA ground pass “Outpost” sessions and new day/night times.  There also is another tagup scheduled at 4:00pm with EVA specialists at TsUP/Moscow to discuss tomorrow’s venture via S-band S/G (space-to-ground).   [As originally planned for last week, after egressing the DC1 at ~5:46pm EDT, the first 30-35 minutes will be spent on preparing the GStM-1 “Strela” cargo boom, mounted on the DC1 hull, and extending it (by manually cranking its pitch and yaw drive handles) towards the FGB, where its end effector (TU) will be attached at a handrail (#1076).  The crane, measuring 3 m retracted and 15 m in fully extended configuration, then serves as handrail and attachment for the crew’s safety tethers until U.S. tethers are secured at the PMA-1 (Pressurized Mating Adapter 1) and later on the S0 truss.  Removal and replacement (R&R) of the failed RPCM, to start at or about 8:29pm, requires prior opening of the door of the portside power assembly SPDA (secondary power distribution assembly).  After the R&R and restoration of external configurations, including retraction and securing of the GStM-1, ingress of the last spacewalker, Padalka (EV1), is expected at ~11:33pm.  The two Mission Control Centers will take turn in leading the spacewalk, with TsUP in charge of Russian segment (RS) operations and egress, translation, return, ingress and any Russian task activities, while MCC-Houston has the lead for detailed ops dealing with the U.S. segment, the RPCM R&R and the power-down of the external Lab DDCU (dc-to-dc converter unit) 2A power source of the RPCM prior to the R&R, and its subsequent power-up.]

Shortly before sleep time tonight, the CDR will once more break out the “Urolux” equipment, setting it up for the Russian biochemical urine test (PZE MO-9), a standard requirement before and after Orlan-suited activities.   [The MO-9 analysis uses the sophisticated in-vitro diagnostic apparatus “Urolux” developed originally for the Mir program.]

Earlier today, Mike Fincke collected and stowed the two FMK (formaldehyde monitoring kit) sampling badges deployed previously (6/11) in the Lab and SM to catch any atmospheric formaldehyde on a sampling substrate for later analysis in JSC lab facilities.

Gennady is scheduled to conduct the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO).

Both crewmembers will perform their daily 2.5-hr. physical exercise program on TVIS, RED exerciser, and VELO cycle with load trainer.

In a change from last week’s plan, during the EVA MCC-Houston will conduct a spin-up test of CMG-2 (control moment gyroscope 2), right after installation and activation of the spare RPCM replacing the failed RPCM S02B-D.  The rotor spin-up is planned for 60 seconds duration.

After the CMG-2 spin-up test, about one hour before the end of the spacewalk, there will also be an attitude control handover test from the Russian segment (RS) to the U.S. segment (USOS) to characterize a disturbance torque observed on the RS during last week’s EVA.

Following the completion of the RPCM R&R, MCC-H will conduct a gimbal bias evaluation on CMG-2, preparatory to introducing the gyro back into the steering law on Friday morning (about 30 hrs after the EVA).  To support the gimbal bias evaluation as well as a Russian GPS (Global Positioning System) test, the ISS will remain in LVLH XVV attitude through 7/2.

U.S. and Russian specialists have worked out a joint plan for post-EVA repressurization of the ISS interior.   [After the EVA, during which cabin pressure is expected to be at ~720 Torr (mmHg), the cabin will first be repressurized with the air remaining in Progress 14P storage, raising total pressure to 730-735 Torr, with a ppO2 (oxygen partial pressure) of ~23%.  To relieve subsequent oxygen production requirements on the Elektron, an additional 19-20 Torr increase from the U.S. N2 HPTs (nitrogen high-pressure tanks) outside the Airlock will then follow.]

Another set of options for Mike’s Saturday Science program this weekend (7/3) was uplinked for his choice.  They are a BCAT-3 (Binary Colloid Alloy Test 3) photo session, test sequences for the CFE (Capillary Flow Experiment), or an EPO (Educational Program Operation) demo of pollution research.

The planned Russian SM 7.02 software update is now scheduled for 7/15-16.  The upload will require inhibiting thrusters for a total of about 15 min.   [The new software includes improvements for ATV (automated transfer vehicle) integration, payload control, and other general upgrades (like software efficiency, operational upgrades, accommodation of interface changes, etc.).]

During the last five days, MCC-H specialists have conducted a plasma characterization test of the ISS external environment using the PCU-2 (Plasma Contactor Unit #2) in Discharge mode (PCU-1 in Standby), which ended today.   [PCU xenon-emission current is the sum of the electron currents collected from the ambient ionosphere by (1) the high-voltage (160V) P6 solar arrays (when in ram position) and (2) via low-voltage induction by the ISS structure moving through the ionosphere.  The PCUs prevent the build-up of potentially hazardous floating point (FP) voltages on conducting ISS structure by emitting the collected electrons back to the ionospheric plasma.  The Russian segment (RS) has been specifically designed to inhibit charge collection.  The current test will help understand the physics of ISS charging and the ability of the RS to inhibit it.  In particular, the data will assist in determining the impact to ISS charging of adding additional SAWs (solar array wings).]

Today’s CEO (Crew Earth Observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Chicago, Illinois (it looked like the weather was clearing with a strong cold front passing.  The crew had a nadir pass over the “Windy City” as ISS approached from NW), Western Equatorial Atlantic (as ISS approached the coast of South America from the NW, the Amazon delta and estuaries were right of track.  Looking left of track for sun glint and internal waves as the Amazon discharge interacts with the open Atlantic. If possible, the crew was to try to include some landmarks in their camera field of view), and Typhoon Tingting (DYNAMIC EVENT:  One of two large typhoon systems active in the western Pacific now, Tingting is the only one the crew was likely to see today.  Looking well right of track, possibly a pan for this Category 3 tropical cyclone).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. & Russian Segment Status  (as of today, 1:00pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.6; ppO2 (mmHg) — 150.5; ppCO2 (mmHg) — 4.7.
  • SM Transfer Compartment:  Pressure (mmHg) — 741; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 755; temperature (deg C) — 20.3.
  • Node:  Pressure (mmHg) — 740.60; temperature (deg C) — 23.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 751.50; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 751.50; temperature (deg C) — 22.6; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3963 kg (8737 lb) as of 6/24/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 8:18am EDT [= epoch]):

  • Mean altitude — 360.5 km
  • Apogee — 364.2 km
  • Perigee — 356.8 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6319 deg
  • Eccentricity — 0.0005451
  • Solar Beta Angle — 25.0 deg
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32025
  • For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.